OpenStudy (anonymous):

What is the y-intercept for the graph of 7x – 3y = –5?

OpenStudy (igreen):

Put it in the form of y = mx + b.

OpenStudy (anonymous):

We need this in point slope form. We must isolate y. 7x - 3y = -5 -3y = -7x - 5 y = -7/-3x -5/-3

OpenStudy (anonymous):

So the y intercept is -5/-3.

OpenStudy (igreen):

7x - 3y = -5 -3y = -7x - 5 y = -7/3x - 5/-3